OpenStudy (vianne):

The price of a car has been reduced from $11,500 to $7,130. What is the percentage decrease of the price of the car?

OpenStudy (vianne):

Please explain your logic as well!

OpenStudy (anonymous):

(7,130/11,500)X100

OpenStudy (vianne):

That'd equal 62% But it's wrong.. :/

OpenStudy (anonymous):

divide and the * by 100

OpenStudy (vianne):

I did

OpenStudy (anonymous):

what are the answer choices?

OpenStudy (anonymous):

43%

OpenStudy (vianne):

There are no answer choices, just a fill in box.

OpenStudy (anonymous):

that is the decrease in the starting price , if you want the reduction % than 100%-62%=38%

OpenStudy (vianne):

What about the increasing %?

OpenStudy (anonymous):

The Price of car decreased , what increase are you referring to

OpenStudy (vianne):

Well, 38% was correct, but I have another question that is the opposite of decrease.. It's: The cost of a jacket increased from $50.00 to $57.00. What is the percentage increase of the cost of the jacket?

OpenStudy (anonymous):

(57,000/50,000 )X100 that is the increase from the starting value , minus from 100% you will get the percentage increase

OpenStudy (anonymous):

11,500 - 7,130 = ? ? / 11,500 = % could this be it?

OpenStudy (anonymous):

its 14%

OpenStudy (vianne):

Thanks, I'll try to do the other questions :3

OpenStudy (vianne):

A company is expanding its building area from 15,500 square feet to 21,390 square feet. What is the percentage increase of the area of the building space? You'd do (15,500/21,390)X100 right?

OpenStudy (anonymous):

(The Second Value / The First Value ) X100 gives the total increase or decrease , if its increase then minus -100% , if its decrease then 100%-value.

OpenStudy (anonymous):

Thats 38%

OpenStudy (vianne):

Oh

OpenStudy (vianne):

The value of a professional basketball player's autograph rose 30% in the last year. It is now worth $325.00. What was it worth a year ago? What about this one?

OpenStudy (anonymous):

Then 325.00 is the new value , Divide By 130% you will get the 1% then multiply By 100 you will get the 100% which is the old autograph.

OpenStudy (anonymous):

269$ i think

OpenStudy (vianne):

$250 325 / .130 = 2500 X 100 = 250000 If you put the decimal it'd be $250.00 :D

OpenStudy (vianne):

Can you help me with this question as well? A pair of shoes usually sells for $55. If the shoes are 20% off, and sales tax is 8%, what is the total price of the shoes, including tax? Insert Logic As Well :)

OpenStudy (anonymous):

Ya , I meant 249.5 , My calc is a bit sensitive.

OpenStudy (anonymous):

Sale tax is abduct able from the ending sale price right?

OpenStudy (anonymous):

so you decrease the 55 by 20% then abduct the taxes which is the 8% from the new price by the same method

OpenStudy (vianne):

Ohk

OpenStudy (anonymous):

(55/100) x20=11 , then the new price is 55-11=44 , then abduct the taxes which is 8% from the ending price (44/100)x8=3.52 , then the final value for the price will be 44-3.52=40.48.

OpenStudy (anonymous):

If the tax is abucted from the initial price then the value will be 39.6 as it will be dec by 28%

OpenStudy (vianne):

That answer isn't correct.. The answer choices for this question is: A $48.40 B $50.49 C $44.55 D $47.52

OpenStudy (anonymous):

its 40.48 , not 48.40 check again.

OpenStudy (anonymous):

w8 , You mean then total price including the tax , so tax is added to the price?

OpenStudy (vianne):

Yep

OpenStudy (anonymous):

it will be 47.52$

OpenStudy (anonymous):

i thought taxes are abducted not added ,sry.

OpenStudy (vianne):

Oh..

OpenStudy (anonymous):

so instead decreasing the tax value you add it to the final price , which is the 8% from the ending price which is 44.

OpenStudy (vianne):

Okay

OpenStudy (anonymous):

(44/100 )x8 +44=47.52

OpenStudy (anonymous):

Thats it for now , this was fun , but i need to go study , bya

OpenStudy (vianne):

Okay.

OpenStudy (vianne):

Bya, I'll fan you! =)